A body is thrown vertically upwards with an initial velocity u reaches maximum height in 6 seconds. The ratio of distance travelled by the body in the first second and seventh second is:
A.)  1 : 1
B.)  11:1
C.)  1 : 2
D.)  1 : 11

Grade:11

1 Answers

Soumya Ranjan Mohanty
6 years ago
Dear student,
for full motion,
v=u-gt
=>0=u-gt
=>u=gt=10*6=60 ms^-1
 
dist travelled in first second=s1= u-(g/2)*(2n-1) = 60-5*1=55
dist travelled in seventh second=s7= u-(g/2)*(2n-1) =60-5*13=-5 (negative as direction is different)
so ratio =s1:|s7|
=55:5
=11:1 (B)
so B is the correct answer.
Thanks,
Hope it helped.
Soumya Ranjan Mohanty.